# Computational Complexity of Physical Information Sufficiency

### v3

@@ 1 Introduction
This is coordinate sufficiency, the formalization of information sufficiency (whether available information determines the optimal action) over a factored state space.
In information-theoretic terms, this is a task-conditional coding problem: the object preserved by projection is the decision boundary $s\mapsto\text{Opt}(s)$, not the full source state. The same object appears in classical source-coding and fidelity viewpoints [57, 59, 14, 64], and in zero-error/confusability structure where exact recoverability depends on graph constraints rather than raw description length [58, 31, 38, 15]. The same separation also appears in model selection: short descriptions and certifiable adequacy are distinct [54, 24, 35, 8].
For physical instantiations, the same tuple has a direct substrate reading: coordinates are measurable state degrees of freedom, actions are physically realizable interventions or transitions, and the optimizer map is the decision boundary induced by the declared objective contract. The physical-to-core encoding discipline in Section 2.7 makes this bridge theorem-level rather than metaphorical.[C:P2.34,2.36,2.28,2.29,2.30,2.31;C2.35;T2.37;R:AR] (L: CT6,CT8,CT4,CT1,DQ548,DQ537,DQ539,DQ541.)
Compact claim-stamp notation.
We use compact provenance stamps of the form $C:\cdots;R:\cdots]$: claim references are grouped by type ($D/P/C/T/L/R$ for Definition/Proposition/Corollary/Theorem/Lemma/Remark), and regime tags are abbreviated (for example ARAR=active declared regime, DCDC=declared contract, CRCR=declared class/regime, QfQf=query-finite, QbQb=query-Boolean). Theorem-strength narrative statements are stamped and, where available, paired with compact Lean handle IDs; unstamped narrative remains explanatory or citation-only context.
SUFFICIENCY-CHECK, MINIMUM-SUFFICIENT-SET, and ANCHOR-SUFFICIENCY are meta-problems: each is parameterized by an arbitrary decision problem $\mathcal{D}$ and asks a structural question about $\mathcal{D}$’s coordinate space. The complexity results therefore apply to every problem with factored state space, not to a specific problem domain.[C:T4.6,4.7,4.9;R:S] (L: CC61,CC36,CC5.)
Remark 1.1 (Universality of the Decision-Problem Tuple).
The formal object of this paper is the decision-problem tuple $\mathcal{D}=(A,S,U)$: actions, factored states, utility. Within this model family, any problem requiring a choice among alternatives given structured information instantiates the tuple: the alternatives are $A$, the structured information is $S=X_{1}\times\cdots\times X_{n}$, and the objective is $U$. Thus the qualifier “decision problem” does not narrow scope inside the declared coordinate model. The meta-problem results below are universal over problems representable with this coordinate structure; the $(A,S,U)$ formalism supplies the language for stating the sufficiency predicate.
This paper characterizes the complexity landscape of these meta-problems within the formal model:
The paper keeps two layers explicit throughout: a substrate-neutral decision core (for complexity classification) and an assumption-typed physical instantiation layer (for transport to concrete systems). This separation prevents conflating representational hardness with physical assumptions while preserving a formal bridge between them.[C:P2.34;T2.37;R:AR] (L: CT6,CT4.)
Section 2.4 fixes the computational model and input encodings used for all complexity claims. Section 2.5 gives the model contract and regime tags used to type every strong claim. Completeness claims are class-internal: complete for the static sufficiency class fixed by C1–C4 and the declared access regimes. Adjacent objective classes that violate the bridge conditions are typed separately (Section 2.5).
Problem Complexity When Tractable
SUFFICIENCY-CHECK coNP-complete Bounded actions (explicit-state), separable utility, tree-structured utility
MINIMUM-SUFFICIENT-SET coNP-complete Same conditions
ANCHOR-SUFFICIENCY $\Sigma_{2}^{\textsf{P}}$-complete Open
The tractable cases are stated with explicit encoding assumptions (Section 2.4). Outside those regimes, the succinct model yields hardness.
1.1 Hardness–Recovery Reading Map
Theorems in this paper are intended to be read as hardness/recovery pairs for the same fixed decision question:
  1. 1.
    Sufficiency checking: [S] coNP-complete in the general succinct regime (Theorem 4.6), paired with polynomial-time recovery under explicit-state and structured-access regimes (Theorem 6.1).
  2. 2.
    Minimal sufficient sets: [S] coNP-complete in general (Theorem 4.7), paired with the collapse criterion and relevance-computation recovery route (Theorems 4.14, 6.1).
  3. 3.
    Anchor sufficiency: $\Sigma_{2}^{\textsf{P}}$-complete in the general anchored formulation (Theorem 4.9); no general polynomial-time recovery condition is established in this model.
(L: CC60, CC36, CC34, CC4, CC71, S2P2.)
An operational criterion follows later from the same chain: once decision-site count exceeds the amortization threshold $n^{*}$, avoiding structural identification is more expensive than paying the one-time centralization cost (Theorem 9.6).[C:T9.6;R:LC] (L: HD3, HD21.)
1.2 Landscape Summary
When is sufficiency checking tractable? We identify three sufficient conditions:
  1. 1.
    Bounded actions ($|A|\leq k$) under explicit-state encoding: with constantly many actions, we enumerate action pairs over the explicit utility table.
  2. 2.
    Separable utility ($U(a,s)=f(a)+g(s)$): The optimal action depends only on $f$, making all coordinates irrelevant to the decision.
  3. 3.
    Tree-structured utility: With explicit local factors over a tree, dynamic programming yields polynomial algorithms in the input length.
Each condition is stated with its encoding assumption. Outside these regimes, the general problem remains coNP-hard (Theorem 4.6). [C:T4.6,6.1;R:E,S] (L: CC61,CC69,CC70,CC71.)
When is it intractable? The general problem is coNP-complete (Theorem 4.6), with a separation between explicit-state tractability and succinct worst-case hardness:
  • In the explicit-state model: SUFFICIENCY-CHECK is solvable in polynomial time in $|S|$ by explicitly computing $\text{Opt}(s)$ for all $s\in S$ and checking all pairs $(s,s^{\prime})$ with equal $I$-projection. In particular, instances with $k^{*}=O(\log|S|)$ are tractable in this model.
  • In the intermediate query-access model: the finite-state core yields an Opt-oracle lower bound $\Omega(|S|)$ (Boolean instantiation: $\Omega(2^{n})$), and the mechanized Boolean interface refinements preserve the obstruction for value-entry and state-batch interfaces, with explicit subproblem-to-full transfer, weighted randomized robustness, and oracle-lattice transfer/strictness theorems (Propositions 5.35.16).
  • In the succinct model: under ETH there exist worst-case instances produced by our strengthened gadget in which the minimal sufficient set has size $\Omega(n)$ (indeed $n$) and SUFFICIENCY-CHECK requires $2^{\Omega(n)}$ time.
The explicit ETH lower bound is still a succinct worst-case statement; Proposition 5.3 provides the finite-state query core and Propositions 5.75.16 provide theorem-level Boolean-interface refinements. [C:P5.3,5.7,5.8,5.11,5.15,5.16;R:Qf,Qb] (L: CC50,L9,L11,L8,L21,L20.)
1.3 Main Theorems
  1. 1.
    Theorem 4.6: SUFFICIENCY-CHECK is coNP-complete via reduction from TAUTOLOGY.
  2. 2.
    Theorem 4.7: MINIMUM-SUFFICIENT-SET is coNP-complete (the $\Sigma_{2}^{\textsf{P}}$ structure collapses).
  3. 3.
    Theorem 4.9: ANCHOR-SUFFICIENCY is $\Sigma_{2}^{\textsf{P}}$-complete via reduction from $\exists\forall$-SAT.
  4. 4.
    Theorem 5.1: Encoding-regime separation: explicit-state polynomial-time (polynomial in $|S|$), and under ETH a succinct worst-case lower bound witnessed by a hard family with $k^{*}=n$.
  5. 5.
    Theorem 6.1: Polynomial algorithms for bounded actions, separable utility, tree structure.
  6. 6.
    Theorem 2.37: physical-system bridge composition (law-objective semantics, behavior-preserving configuration reduction, and required-work lower bound) is proved as one bundled mechanized statement under declared encoding assumptions.
  7. 7.
    Propositions 2.282.31: the interface-time model is discrete, each decision event advances time by one unit, run-length equals elapsed interface time, and the unit-time law is substrate-invariant under the declared interface-consistency condition.
[C:T4.6,4.7,4.9,5.1,6.1,2.37;P2.28,2.29,2.30,2.31;R:E,S,S+ETH,AR] (L: CC61,CC36,CC5,CC16,CC69,CC70,CC71,CT4,DQ548,DQ537,DQ539,DQ541.)
1.4 Claim-Typing Theorems
  1. 1.
    Proposition 3.16: report admissibility is typed by certificate class ($\mathsf{ABSTAIN}$, $\mathsf{EXACT}$, $\mathsf{EPSILON}(\varepsilon)$).
  2. 2.
    Proposition 3.18: first-class evidence-object existence is equivalent to typed report admissibility.
  3. 3.
    Propositions 2.5 and 2.6: the $I=\emptyset$ query is exactly universal constancy of the decision boundary, and non-sufficiency is exactly existence of a counterexample pair.
  4. 4.
    Proposition 3.21: exact admissibility is equivalent to exact evidence existence.
  5. 5.
    Proposition 3.23 and Corollary 3.25: certified confidence is evidence-gated; for exact reports in no-competence regimes, certified confidence must be zero.
  6. 6.
    Proposition 3.24: under signal consistency, no evidence implies zero certified confidence.
  7. 7.
    Proposition 3.29 and Proposition 3.30: abstain may carry tentative guesses and self-reflected confidence, but self-confidence alone does not certify exactness.
  8. 8.
    Propositions 3.26, 3.27, 3.28: scalar step-count witnesses are always exact/falsifiable, while completion percentages are meaningful only under explicit declared bounds.
  9. 9.
    Theorem 3.56: outside declared carve-outs, admissible exact-claim policy requires explicit assumptions.
  10. 10.
    Proposition 3.61: under non-collapse, integrity and exact competence cannot both hold universally on hardness-blocked declared regimes.
  11. 11.
    Proposition 3.62 and Corollary 3.63: schema-level consequence for declared class/regime objects and exact-report inadmissibility.
[C:T3.56;P3.16,3.18,2.5,2.6,3.21,3.23,3.24,3.29,3.30,3.26,3.27,3.28,3.61,3.62;C3.25,3.63;R:DC,AR] (L: CC75,IC2,DP6,DP7,IC35,IC38,IC40,IC46,IC47,IC42,IC44,IC45,CC30,CC15,CC38.)
1.5 Regime Separation Theorems
  1. 1.
    Theorem 5.1: same-question separation between explicit-state and succinct regimes (with [S+ETH] hard-family lower bound).
  2. 2.
    Proposition 5.3: finite-state Opt-oracle lower bound in [Q_fin].
  3. 3.
    Propositions 5.75.16: Boolean-interface refinements and transfer/strictness in [Q_bool].
[C:T5.1;P5.3,5.7,5.11,5.8,5.15,5.16;R:E,S+ETH,Qf,Qb] (L: CC16,CC50,L9,L8,L11,L21,L20,L23.)
1.6 Machine-Checked Proofs
The reduction constructions and key equivalence theorems are machine-checked in Lean 4 [16, 17, 67, 21, 32] (13391 lines, 597 theorem/lemma statements). The formalization verifies that the TAUTOLOGY reduction correctly maps tautologies to sufficient coordinate sets. Complexity class membership (coNP-completeness, $\Sigma_{2}^{P}$-completeness) follows by composition with standard complexity-theoretic results.
What is new.
We contribute (i) formal definitions of decision-theoretic sufficiency in Lean; (ii) machine-checked proofs of reduction correctness for the TAUTOLOGY and $\exists\forall$-SAT reductions; (iii) a complete complexity landscape for coordinate sufficiency with explicit encoding assumptions; and (iv) a formal separation between structural complexity and representational hardness used to derive theorem-indexed engineering corollaries, including theorem-level typed coverage/completeness closures for the declared static class (Theorems 2.38, 3.4). Prior work establishes hardness informally; we formalize the constructions and their regime-typed consequences. [C:T2.38,3.4,4.6,4.9;R:CR,S] (L: CC14,CC76,CC61,CC5.)
@@ 1.7 Paper Structure
[…]
Section 2: decision-problem foundations and encoding model. Section 3: structural vs representational hardness; integrity vs competence. Section 4: hardness proofs. Section 5: regime separation. Section 6: tractable cases. Section 7: regime-conditional corollaries. Section 8: dominance theorems for hardness placement. Section 9: conservation law for decision-relevant coordinates. Section 10: related work. Appendix A: Lean listings.
Sections 79 are not applied commentary. Every claim in those sections is a formal corollary, regime-typed by the framework of Section 3 and indexed to a machine-checked theorem handle. They complete the complexity landscape by characterizing when exact minimization is rational and when it is not: the rationality matrix (Proposition 3.60), the dominance and conservation theorems (Theorems 8.16, 9.3), and the amortization threshold (Theorem 9.6) are all mechanized in Lean. [C:P3.60;T8.16,9.3,9.6;R:AR,LC] (L: IC27,HD1,HD4,HD3.)
@@ 2 Formal Foundations
We formalize decision problems with coordinate structure, sufficiency of coordinate sets, and the decision quotient, drawing on classical decision theory [55, 52]. The core definitions in this section are substrate-neutral: they specify the decision relation independently of implementation medium. Physical interpretation is introduced later through explicit encoding and assumption-transfer maps (Section 2.7).[C:D2.32;P2.34;T2.37;R:AR] (L: CT6,CT4.)
@@ 2.4 Computational Model and Input Encoding
We fix the computational model used by the complexity claims. In physical deployments, these encodings correspond to different access conditions on the same underlying system: explicit tables correspond to fully exposed state/utility structure, succinct encodings correspond to compressed generative descriptions, and query regimes correspond to constrained measurement interfaces.[C:T5.1;P5.3,2.34;R:E,S+ETH,Qf,AR] (L: CC16,CC50,CT6.)
@@
[C:P2.22;R:DM] Let $\mathrm{repr}:\mathcal{L}\to\mathcal{C}$ be any representation map from logical possibilities to code states. If $|\mathcal{C}|<|\mathcal{L}|$, then there exist distinct logical states with the same representation:
@@
This is the finite-cardinality collision statement underlying under-resolved exact-decision claims. (L: DQ582, DQ583.)
@@ 2.5 Model Contract and Regime Tags
In this paper, normative modal statements denote contract-conditional necessity: once an agent declares an objective/regime/assumption contract, admissible policy is fixed by typed admissibility and integrity under that contract. No unconditional external moral prescription is asserted.
@@ 2.6 Discrete-Time Event Semantics
For any substrate model that is consistent with the declared interface transition rule, each one-step substrate evolution realizes one interface decision event and advances interface time by one unit. The time update law is invariant under substrate kind. (L: DQ542,DQ541,DQ549.)
@@ 2.7 Physical-System Encoding and Theorem Transport
[C:P2.33;R:AR] Fix a declared noisy physical encoding interface and assume a Heisenberg-strong binding witness: one physical instance is compatible with two distinct interface states whose encoded optimal-action sets differ. Then there exists a core decision problem with non-constant optimizer map. Equivalently, under this declared physical ambiguity assumption, decision nontriviality is derivable at the core level. (L: DQ584, DQ585, DQ586. )
Proposition 2.34 (Conditional Physical Claim Transport).

### v4

# Computational Complexity of Physical Counting

@@ 1.4 Our Formal Tools: Coordinate Sufficiency
We prove complexity classifications for three meta-problems:
  • SUFFICIENCY-CHECK: Is $I$ sufficient? coNP-complete
  • MINIMUM-SUFFICIENT-SET: Find minimal $I$. coNP-complete
  • ANCHOR-SUFFICIENCY: Exists sufficient $I$ with $|I|\leq k$? $\Sigma_{2}^{P}$-complete
Stochastic variants are PP-complete. Sequential variants are PSPACE-complete. Strict hierarchy: static $\subset$ stochastic $\subset$ sequential.
1.5 Machine Verification
All results are machine-checked in Lean 4:
  • 24,347 lines of Lean code
  • 1,082 theorems
  • 0 sorry placeholders
  • 0 hidden axioms
Bayesian optimality: proved from $\log x\leq x-1$ alone. Thermodynamic consequences: conditional on Landauer. Complexity results: hypotheses as explicit theorem parameters. Machine-generated assumption ledger records all dependencies.
@@ 1.6 Paper Structure
Section 2: formal definitions (decision problems, sufficiency, DQ). Section 5: physical grounding (Counting Gap, three laws). Section 4: hardness proofs (coNP, $\Sigma_{2}^{P}$, PP, PSPACE). Section 9: regime separation. Section 10: six tractable subcases. Sections 1113: engineering corollaries. Appendix: Lean listings.
Claim-stamp notation.
Claims are stamped with theorem references $[T:n;P:m]$ and Lean handles ().
Regime Question Placement
Static SUFFICIENCY / MINIMUM / ANCHOR coNP / coNP / $\Sigma_{2}^{P}$
Stochastic STOCHASTIC-SUFFICIENCY $P$-complete
Sequential SEQUENTIAL-SUFFICIENCY PSPACE-complete
(L: QT1, QT2, QT3, WD1, WD2, WD3, BC1, BC2, BC3, BC4, BC5, CCC6, CCC5, CCC1, DC9, DC10, DC1, DC2, FN14, MN5, FI7, W1, W2, W3, W4, RD1, RD2, RD3, RS1, RS2, RS3, RS4, RS5, TUR1, TUR2, TUR5, TUR6.)
@@ 2 Formal Foundations
We formalize decision problems with coordinate structure, sufficiency of coordinate sets, and the decision quotient, drawing on classical decision theory [73, 69]. The core definitions in this section are substrate-neutral: they specify the decision relation independently of implementation medium. Physical interpretation is introduced later through explicit encoding and assumption-transfer maps (Section 2.7).[C:D2.38;P2.42;T2.45;R:AR] (L: CT6, CT4.) Physical-state semantics are typed by a realizable state set: any state marked physical has an instance witness, and state-level claims transport over all physical states without substrate-specific axioms. (L: PS1, PS2, PS3, PS4.) Observer classes induce effective quotient state spaces on the same physical domain; witness pairs certify observer-relative collapses/separations. (L: OR3, OR4, OR5, OR9.) Two recurrent circuits coupled by a shared channel node are formalized with clamp-observed state, one-bit decision-event lower bounds, and typed YES/NO/ABSTAIN report projection into the claim-admissibility layer. (L: CV4, CV5, CV7, CV17, CV18, CV11, CV12, CV13, CV15, CV16.) Under separating physical parameterizations, logic/time scaffolds do not determine a unique timed law/objective; deadline forcing yields forced action with positive Landauer/NV lower bounds, and with uncertainty yields nondegenerate belief. (L: CF2, CF4, CF6, CF8, CF9, CF7.) Physical-collapse transfer is isolated as an assumption-indexed chain at requirement and canonical levels: declared collapse-map schema implies no physical collapse and yields the canonical P=NPP{=}NP physical-impossibility specialization. The map is also derived from an explicit SAT witness plus reduction-energy transport, and collapse is bounded by the sharp disjunction (unbounded budget or zero bit-cost or failed exponential lower bound) under bounded-budget/positive-cost/exponential-lower-bound premises; deterministic/probabilistic/sequential variants preserve the same form. (L: PH11, PH12, PH13, PH14, PH15, PH16, PH17, PH18, PH19, PH20, PH21, PH22, PH23, PH24, PH25, PH26, PH27, PH28, PH29, PH30, PH31, PH32, PH33.) A layered graph module composes cycle witnesses, observer surprisal-plus-novelty nontriviality, entropy-normalized DQ, forced-belief/Bayes-update bridge, and thermodynamic/material lower-bound witnesses. (L: GN2, GN4, GN5, GN6, GN7, GN8, GN9, GN10, GN11, GN12, GN13.) It also fixes the exact-claim admissibility boundary: exact reports require a certified-bit gap over raw payload, hardness-blocked exact claims collapse to raw-only accounting, and integrity/competence cannot both hold universally under hardness-blocked resources.[C:T3.47;P3.61;C3.49;R:AR] (L: CC18, CC30, CC21.)
@@ 2.1 Decision Problems with Coordinate Structure
Formally: this subsection fixes the core tuple, projection, optimizer, and sufficiency/relevance predicates used downstream.
@@ 2.2 Sufficiency and Relevance
Informally: all later results use this same sufficiency definition.
@@ 2.4 Computational Model and Input Encoding
Formally: this subsection defines encoding/access regimes, cost measures, and regime-specific complexity interpretation.
We fix the computational model used by the complexity claims. In physical deployments, these encodings correspond to different access conditions on the same underlying system: explicit tables correspond to fully exposed state/utility structure, succinct encodings correspond to compressed generative descriptions, and query regimes correspond to constrained measurement interfaces.[C:T9.1;P9.3, 2.42;R:E, S+ETH, Qf, AR] (L: CC16, CC50, CT6.)
@@
Definition 2.22 (Anchor Query Object).
An anchor query object is a pair
$q=(\mathcal{D},I)$
where $\mathcal{D}$ is a decision problem and $I$ is a fixed coordinate set. Its truth predicate is
$\mathrm{AnchorTrue}(q)\iff\exists s_{0}\in S\ \forall s\in S,\ s_{I}=s_{0,I}\Rightarrow\text{Opt}(s)=\text{Opt}(s_{0}).$
Definition 2.23 (Pose Operation for Anchor Queries).
The pose operation is the constructor
$\mathrm{poseAnchor}:\ (\mathcal{D},I)\mapsto q.$
Posing is an operation; $q$ is the formal object.
Proposition 2.24 (Posing Produces a Formal Query Object).
For every $(\mathcal{D},I)$, $\mathrm{poseAnchor}(\mathcal{D},I)$ is a well-typed anchor query object with preserved components, and its truth predicate is exactly the anchored $\exists\forall$ condition in Definition 2.22. [C:D2.22, 2.23;T4.9;R:S] (L: AQ1, AQ2, AQ3.)
Proposition 2.25 (Typed Exact Claims on Posed Anchor Queries).
For posed anchor queries under the typed claim discipline:
$\text{Exact admissible}\iff\text{competent on the declared regime},\qquad\neg\text{competent}\Rightarrow\neg\text{exact admissible}.$
Under solver integrity, checker-accepted exact-true verdicts are semantically sound on the posed anchor query object. [C:P3.16, 3.58;C3.50;R:AR] (L: AQ4, AQ5, AQ6, AQ7.)
Corollary 2.26 (Certified-Confidence Gate for Posed Anchor Queries).
Under signal consistency, positive certified confidence implies typed admissibility of the posed anchor-query report. [C:D3.20;P3.23;R:AR] (L: AQ8.)
@@
[C:P2.27;R:DM] Let $\mathrm{repr}:\mathcal{L}\to\mathcal{C}$ be any representation map from logical possibilities to code states. If $|\mathcal{C}|<|\mathcal{L}|$, then there exist distinct logical states with the same representation:
@@
This is the finite-cardinality collision statement underlying under-resolved exact-decision claims. (L: PI6, PI7.)
Theorem 2.28 (Physical Incompleteness of Instantiation).
[C:T2.28;R:AR] Let $\mathcal{P}$ be a finite set of physical states and $\mathcal{L}$ a finite set of logical possibilities, with an instantiation map $\iota:\mathcal{P}\to\mathcal{L}$. If $|\mathcal{P}|<|\mathcal{L}|$, then:
  1. 1.
    $\iota$ is not surjective;
  2. 2.
    there exists a logical possibility $\ell\in\mathcal{L}$ that is not physically instantiated: $\ell\notin\mathrm{range}(\iota)$.
Under explicit bounds: if $|\mathcal{P}|\leq M$ and $|\mathcal{L}|\geq L$ with $M<L$, the same conclusion holds. (L: PI3, PI4, PI5.)
@@
Informally: the same question can be easy or hard depending on representation and access, with explicit transfer requirements.
@@ 2.5 Model Contract and Regime Tags
In this paper, normative modal statements denote contract-conditional necessity: once a physical-system model declares an objective/regime/assumption contract, admissible reports and interventions are fixed by typed admissibility and integrity under that contract. No unconditional external moral prescription is asserted.
@@ 2.6 Discrete-Time Event Semantics
For any substrate model that is consistent with the declared interface transition rule, each one-step substrate evolution realizes one interface decision event and advances interface time by one unit. The time update law is invariant under substrate kind. (L: DT22, DT23, DT24.)
@@ 2.7 Physical-System Encoding and Theorem Transport
Formally: this subsection defines physical-to-core encoding and proves theorem transport via declared assumption maps.
@@
Definition 2.39 (Physical Scope Gate).
A claim is in physical scope only if it declares:
  1. 1.
    a physical state class with observable coordinates,
  2. 2.
    a physical-to-core encoding map $E:\mathcal{P}\to\mathcal{D}$,
  3. 3.
    calibrated measurement units for reported quantities (for example bits, joules, and $kT$-scaled bounds),
  4. 4.
    a declared objective contract and admissible intervention class,
  5. 5.
    an explicit physical-to-core assumption transfer map.
Claims missing any item are untyped for theorem transport in this framework.
Proposition 2.40 (Typed Physical Transport Requirement).
Direct transfer of core theorems is licensed only through a declared physical scope gate instance (Definition 2.39) and a corresponding physical-to-core encoding (Definition 2.38). Absent that typed physical instance, transport is blocked by construction. PS3PS4
Proof.
By Definition 2.39, theorem transport is only defined for claims with explicit physical state, objective, intervention, encoding, and assumption-transfer declarations. Therefore untyped targets have no admissible transport map in this framework. Proposition 2.42 then applies only to typed instances.  $\blacksquare$
This scope gate is a typing rule, not a rhetorical caveat: the paper’s transport theorems apply only after explicit encoding and assumption transfer are declared.[C:D2.38, 2.39;P2.40, 2.42;R:AR]
@@
[C:P2.41;R:AR] Fix a declared noisy physical encoding interface and assume a Heisenberg-strong binding witness: one physical instance is compatible with two distinct interface states whose encoded optimal-action sets differ. Then there exists a core decision problem with non-constant optimizer map. Equivalently, under this declared physical ambiguity assumption, decision nontriviality is derivable at the core level. (L: HS3, HS5, HS6.)
Proposition 2.42 (Physical Claim Transport).
@@
Informally: physical conclusions apply when the stated physical assumptions line up with the core theorem assumptions.